OpenStudy (anonymous):

If logV = log4 - log3 +logπ +3logr, then write a formula for V in terms of π and r

OpenStudy (anonymous):

log4/3 πr³Deleting log

OpenStudy (anonymous):

wait im wrong somewhere

OpenStudy (anonymous):

wat is the variable here

OpenStudy (anonymous):

combine the right hand side in to a single log using \[\log(a)+\log(b)=\log(ab)\] \[\log(a)-\log(b)=\log(\frac{a}{b})\] \[n\log(a)=\log(a^n)\]

OpenStudy (anonymous):

\[\log_{} V=\log_{} 4/3+\log_{} r^3*\pi\] \[\log_{} V=\log_{} ([4r^3*\pi]/3)\] \[V=4r^3*\pi/3\]

OpenStudy (anonymous):

\[\log(4)+\log(\pi)+3\log(r)-\log(3)\] \[=\log(4)+\log(\pi)+\log(r^3)-\log(3)\] \[=\log(\frac{4\pi r^3}{3})\] if \[\log(V)=\log(\frac{4\pi r^3}{3})\] then \[V=\frac{4\pi r^3}{3}\]